OpenStudy (anonymous):

A card is drawn from a standard deck of cards. Determine whether the events are mutually exclusive or inclusive. Then, find the probability. P(ace or club) a. inclusive, 4/13 b. mutually exclusive, 17/52 c. mutually exclusive, 4/13 d. inclusive, 17/52 Can you explain this to me step by step ? I actually want to know how to do it.

OpenStudy (nincompoop):

post all of your homework questions in one thread so people can feast on them with response. i don't think you're even trying to solve any of them by yourself.

OpenStudy (anonymous):

i am trying, the ones i'm posting up are the ones i don't understand at all.

OpenStudy (anonymous):

you know the formula for probablilty (what you want) / (total outcomes) yes?

OpenStudy (anonymous):

no i don't even know how to start the problem

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

probablilty is calculating the fraction of something you want to happen. in your case you want to draw either a ace or a club from a deck of cards. so start simpler, do you know the probablity of pulling a 4 from a deck of cards?

OpenStudy (anonymous):

so that's 4/52 which will be 1/13?

OpenStudy (anonymous):

yes good... what you have is a mutually exclusive event because you can have BOTH an ace and a club at the same time... so there's a special formula for that

OpenStudy (anonymous):

P(A or B) = P(A) + P(B) - P(A and B)

OpenStudy (anonymous):

so we'll say that there are 52 cards in the deck and that will be our denominator, we'll focus on the numerators. How many aces are in a deck of cards?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

ok so P(A) = 4/52... how many clubs are in a deck of cards?

OpenStudy (anonymous):

1/13?

OpenStudy (anonymous):

no, there's 13 clubs in a deck of cards so P(B) = 13/52

OpenStudy (anonymous):

now here's the tricky part of the problem is that there is a CHANCE of getting both AN ACE and A CLUB at the same time. that's what makes it mutually exclusive so you have to subtract this chance away so you don't count it twice. so P(A) = 4/52 P(B) = 13/52 and P(A and B) = 1/52. So what is P(A) + P(B) - P(A or B)?

OpenStudy (anonymous):

oops should be what is P(A) + P(B) - P(A and B)

OpenStudy (anonymous):

P(4/52) + P(13/52) - P(4/52 and 13/52) P(17/52) - P(4/52 and 13/52) so on the P(A and B) part do i add them together?

OpenStudy (anonymous):

no P(A and B) is the probablity of getting BOTH an ace and a club. how many aces of clubs are in a deck of cards?

OpenStudy (anonymous):

13/52

OpenStudy (anonymous):

remind me not to play cards with you :) there's 13 clubs in a deck of cards, how many of those clubs are aces also?

OpenStudy (anonymous):

there's 13 clubs but 4 of them are aces

OpenStudy (anonymous):

no there is only 1 ace of clubs elsa

OpenStudy (anonymous):

there are 4 aces TOTAL in the deck, but not all of the aces are clubs

OpenStudy (anonymous):

the probablilty of drawing an ace and a club at the same time is 1/52

OpenStudy (anonymous):

P(A) = probality of drawing an ace = 4/52 P(B) = probability of drawing a club = 13/52 P(A and B) = probability of drawing both at the same time = 1/13 P(A or B) = P(A) + P(B) - P(A or B) = (4 + 13 - 1) / 52

OpenStudy (anonymous):

grr last line should read: P(A or B) = P(A) + P(B) - P(A and B) = (4 + 13 - 1) / 52